OpenStudy (debpriya):

Why is the determinant of x'x(x transpose x) 0 ?

OpenStudy (tkhunny):

multiplied by zero?

OpenStudy (debpriya):

No no the determinant of x'x is 0 ...my question is how ?

OpenStudy (eliesaab):

\[ x=\left( \begin{array}{cc} 1 & 2 \\ 3 & 5 \\ \end{array} \right) \]

OpenStudy (eliesaab):

\[ x'=\left( \begin{array}{cc} 1 & 3 \\ 2 & 5 \\ \end{array} \right) \]

OpenStudy (eliesaab):

\[ x x'=\left( \begin{array}{cc} 5 & 13 \\ 13 & 34 \\ \end{array} \right)\]

OpenStudy (eliesaab):

So Det(x x')=1, so it is not zero in general

OpenStudy (eliesaab):

Also Det[ x'x]=1